Which equation would make this system consistent and dependent?
2/3y=7x-1/2
A. 21x=y+3/2 B. 42x-4y=3
C. y = 21/2x-3/2 D. 21x=2y-3
When I put the equations in slope intercept form C comes out as my answer (having the same line). The book is telling me that the correct answer is B. Why is B the answer?

Answer by Boreal(15235)

A. 2/3y=7x-1/2 ; y=(21/2)x-(3/4)
B. 42x-4y=3 ; 4y=42x-3; y=(21/2)x-(3/4)
C. y = 21/2x-3/2 ; the slopes are the same, but the intercept is different. Those lines are parallel but not the same line
D. 21x=2y-3 ; y=(21/2)x+(3/2)